Olongapo Sports Corporation distributes two premium golf balls—Flight Dynamic and Sure Shot. Monthly sales and the contribution margin ratios for the two products follow:

​​​​​​​Product Flight Sure Dynamic Shot $ 690,000 $ 310,000 69% 78% Sales CM ratio Total $ 1,000,000 ? Fixed expenses total $578,500
Answer #1
Required 1:
FIGHT DYNAMIC SURE SHOT TOTAL COMPANY
AMOUNT % AMOUNT % AMOUNT %
SALES 690000 100 310000 100 1000000 100
VARIABLE EXPENSES 213900 31 68200 22 282100 28
CONTRIBUTION MARGIN 476100 69 241800 78 717900 72
FIXED EXPENSES 578500
NET OPERATING INCOME 139400
*CONTRIBUTION MARGIN(CM Ratio)% (476100+241800)/1000000 = 72%
Required 2:
Break even point = fixed cost/total CM Ratio
i.e, = 578500/72%
       = 578500/.72
      = $803472
Required 3:

The additional contribution margin from the additional sales is computed as follows:$55,000 × 72% CM ratio = $39,600.         

Assuming no change in fixed expenses, all of this additional contribution margin of $39,600 should drop to the bottom line as increased net operating income.

This answer assumes no change in selling prices, variable costs per unit, fixed expense, or sales mix